OpenStudy (anonymous):

Can someone help to graph g(X)=abs(x+2)-abs(x-3)+1 , by turing it into a piecewise

OpenStudy (anonymous):

by looking at some pdfs, i was able to see that we could first work with the term abs(x+2), we woudl get : g(x)={ - (x+2)-|x-3|+1,x+2<0 {(x+2)-|x-3|+1,x+2>=0

OpenStudy (anonymous):

then that would simplify to: g(x)={-x-1-|x-3|, if x<-2 {x+3-|x-3|, if x>=-2

OpenStudy (anonymous):

then we work with the term abs(x-3), which goes like this: |x-3|={-(x-3), if x-3<0 {(x-3), if x-3>=0 which simplifies too: |x-3|={-x+3, if x<3 {x-3, if x>=3

OpenStudy (anonymous):

now here the author, i feel kinda of just went straigt to the solution, without really expaling. So what i did was try and figure it out by my own terms, i said, well we have a number line broken up by 2 points:

OpenStudy (mathmate):

Start with f(x)=abs(x), which is a V-shaped graph at 45 degrees with the axes. Using the rules of transformations, f(x+2) means shifting f(x) to the left by 2 units, and similarly f(x-3) means shifting f(x) to the right by 3 units. Also, -f(x) = inverting the graph, i.e. flipping the graph about the x-axis. Combining all these, we have to combine the three together.